OpenStudy (anonymous):

How to get t from h(t) = -4t^2 + 32t using factored form?

OpenStudy (anonymous):

Factored: t(-4t+32) = 0 t = 0, 8

OpenStudy (anonymous):

I am still so confused. h(t) = -4t^2 + 32t 0 = -4t(t-8) and then what's the next step?

OpenStudy (anonymous):

Well with it factored in that form, in order for the equality to be true (that is, equal to zero), either -4t is 0 or t-8 is zero. So t equals 0 or 8

OpenStudy (anonymous):

You take each factored expression and equate them separately to zero. -4t = 0 and t - 8 = 0

OpenStudy (anonymous):

why do I do that?

OpenStudy (anonymous):

You are trying to find which values of t make the polynomial zero.

OpenStudy (anonymous):

ohh, i see. Thank you!

OpenStudy (anonymous):

If \[ab = 0\] then either a is 0 or b is 0. In this case, a=-4t and b=t-8, so you set a and b to 0 separately and solve them.

OpenStudy (anonymous):

ohhh.. I get it. thank you too!